Differential Equations Class, please give general solution

1 -1 2 8. x(t) 4 -1 ) x(t) 3 2 1 9. I(t) = 3 4 -2 -1 î) (0)
Answer #1

3 2 2 17 Giver x(x): 1 2 3 4) This is of th form X(t) = AX (4) where Lel an eigenvalue of A then del (A -AI) to 1 Abe 1-2 -1R2-R3 - 1 2 2 m3 - 1 -2 R3 R3 -2 R2 0 O -1 4 R, R, + R3 lio 0 42 -4 ng OI -4. 23 nit ng so 2, = -83 ng - 423=0 h2 = 423 ชา) uX(4): .84 1-c, etc, ezt 4 C, et + C282t + 2 Cg est Cze (ciet + czë where (,, (2463 are constant - 26 + Cze Answer, 3/4)=(-:)The characteristic equation of a 8? - 27 + 5 = 8= 2 + 4-20 2 I 41 2 2 > this is where * = 1121 of the form 8 &t ß q=1 4 B=2 rTherefore X () = ( 27 (1) is the required half)) general solution where brill) f delt) is given by equation ③ to respectively